OpenStudy (anonymous):

What is the solution of the system? Use substitution. y = x + 6 y = 2x A. (6, 12) B. (–12, –6) C. (–6, –12) D. (2, 4)

OpenStudy (kymber):

Change around one equation to equal x & then plug it into the other one

OpenStudy (anonymous):

you can get y=x+6y (1) ->x+5y=0 (3) y=2x (2) let (3)into (2)

OpenStudy (anonymous):

i think there are something wrong of the equations

OpenStudy (anonymous):

i think it's B